# Reconfigurable Stochastic Neurons Based on Strain Engineered Low Barrier Nanomagnets

### v1

@@ 2 Strain effects on energy barrier
The steady-state in-plane potential energy in a strained nanomagnet with in-plane anisotropy depends on the magnetization orientation as [5]
@@
We carried out Landau-Lifshitz-Gilbert (LLG) simulations of the magnetization dynamics in an LBM at room temperature under different strains to see how the magnetization fluctuation behaves. The LBM we studied is an elliptical Co nanomagnet of major axis 100 nm, minor axis 99 nm and thickness 5 nm. A nanomagnet of these dimensions are likely to be monodomain and hence the macrospin approximation holds. The saturation magnetization = 10 A/m, the magnetostriction coefficient = -35 ppm and the Gilbert damping coefficient = 0.01 correspond to a Co nanomagnet. The coupled LLG equations governing the temporal evolutions of the scalar components of the magnetization were solved with finite difference method [15, 16] with a time step of 0.1 ps. We assumed positive (tensile) stress since Co has negative magnetostriction. The initial condition was that the magnetization was aligned along the minor axis of the nanomagnet.
@@
where is the Gilbert damping factor of the nanomagnet material, is the gyromagnetic factor (a constant), is the -th component of the magnetization at time , and is the -th component of the effective magnetic field experienced by the nanomagnet at time .
@@
where = 1 - - and with () being three uncorrelated Gaussians of zero mean and unit standard deviation, is the nanomagnet volume and is the attempt period which is the time step of the simulation.
@@ 3.1 Energy cost of reconfiguration
We can make an order estimate of the energy cost of reconfiguration. This is the quantity , where is the capacitance of the two gate pads in Fig. 3 and is the gate voltage needed to generate the required stress. We see from Fig. 4 that 6.5 MPa of stress is enough to reconfigure a BSN into an ASN. The gate voltage needed to generate a given stress is given by , where is the diagonal element of the piezoelectric tensor, is the Young’s modulus of the nanomagnet material and is the piezoelectric layer thickness. We will assume that the piezoelectric is PMN-PT whose reported value is 2500 pC/N [17] and that = 300 nm. The Young’s modulus of Co is 209 GPa. This will make the gate voltage needed to generate 6.5 MPa of stress = 3.6 mV. The capacitance of the two gate pads is = , where is the dielectric constant of PMN-PT = 40008.85410 F/m [18] and is the area of the gate pads = 100 nm 100 nm. This makes = 2.4 fF. Hence the energy cost of reconfiguration is only 8.510 Joules, which is miniscule. The gate voltage of 3.6 mV is above the noise voltage at room temperature, which is = 1.3 mV ( = thermal energy). If more noise resilience is desired, one can increase the gate voltage beyond 3.6 mV to obtain the desired noise margin, while still dissipating negligible energy to reconfigure.

### v2

@@ 2 Strain effects on energy barrier
The steady-state in-plane potential energy in a strained elliptical nanomagnet with in-plane anisotropy depends on the magnetization orientation as [5]
@@
We carried out Landau-Lifshitz-Gilbert (LLG) simulations of the magnetization dynamics in an LBM at room temperature under different strains to see how the magnetization fluctuation behaves. The LBM we studied is an elliptical Co nanomagnet of major axis 100 nm, minor axis 99 nm and thickness 5 nm. A nanomagnet of these dimensions are likely to be monodomain and hence the macrospin approximation holds. The saturation magnetization = 106 A/m, the magnetostriction coefficient = -35 ppm and the Gilbert damping coefficient = 0.01 correspond to a Co nanomagnet. The coupled LLG equations governing the temporal evolutions of the scalar components of the magnetization were solved with finite difference method [15, 16] with a time step of 0.1 ps. We assumed positive (tensile) uniaxial strain applied along the major axis of the nanomagnet since Co has negative magnetostriction. This will depress the energy barrier within the nanomagnet. The initial condition was that the magnetization was aligned close to the major axis of the nanomagnet.
@@
where is the Gilbert damping factor of the nanomagnet material, is the gyromagnetic factor (a constant), is the -th component of the magnetization at time , and is the -th component of the effective magnetic field experienced by the nanomagnet at time . The major axis of the nanomagnet is along the y-direction and the minor axis is along the x-direction.
@@
where = 1 - - and with () being three uncorrelated Gaussians of zero mean and unit standard deviation, is the nanomagnet volume, is the Young’s modulus of the nanomagnet, is the magnitude of the strain, and is the attempt period which is the time step of the simulation.
@@ 3.1 Energy cost of reconfiguration
We can make an order estimate of the energy cost of reconfiguration. This is the quantity , where is the capacitance of the two gate pads in Fig. 3 and is the gate voltage needed to generate the required stress. We see from Fig. 4 that 6.5 MPa of stress is enough to reconfigure a BSN into an ASN. The gate voltage needed to generate a given stress is given by , where is the diagonal element of the piezoelectric tensor and is the piezoelectric layer thickness. We will assume that the piezoelectric is PMN-PT whose reported value is 2500 pC/N [17] and that = 300 nm. The Young’s modulus of Co is 209 GPa. This will make the gate voltage needed to generate 6.5 MPa of stress = 3.6 mV. The capacitance of the two gate pads is = , where is the dielectric constant of PMN-PT = 40008.85410-12 F/m [18] and is the area of the gate pads = 100 nm 100 nm. This makes = 2.4 fF. Hence the energy cost of reconfiguration is only 8.510-20 Joules, which is miniscule. The gate voltage of 3.6 mV is above the noise voltage at room temperature, which is = 1.3 mV ( = thermal energy). If more noise resilience is desired, one can increase the gate voltage beyond 3.6 mV to obtain the desired noise margin, while still dissipating negligible energy to reconfigure.